Kayyy4oe:

Each Friday, the school prints 400 copies of the school newsletter. The equation c = 400w models the relationship between the number of weeks and the total number of copies of the newsletters printed. What is true of the graph of this scenario? A viable point on the graph is A.(-2, -800) B.(2.5, 1,000) C.(6, 1,800) D.(8, 3,200)

Kayyy4oe:

I need help

snowflake0531:

do you think it's right or wrong? or are you 100% sure that you 100% dk

Kayyy4oe:

I need help on the problem..

snowflake0531:

Well, it's asking your opinion, whether you think that the equation c=100w is correct so, do you think it's right or wrong, or 100% dk

snowflake0531:

or, in a different way, adding an explanation, the question explicitly wrote 400 copies of it per Friday

Kayyy4oe:

Don’t know because I need help on the problem

snowflake0531:

^^^ 400 copies per Friday, and cuz they only print it on Friday's, it means they print it 400 copies per week *sorry, you're new here, but there's a rule that i can't give direct answer, which in this case is a bit hard lol

snowflake0531:

anyways it means that it's 400 per week so 400*w = the total number of newspapers per w weeks

snowflake0531:

so, it's correct @Tranquility i explained it first? does it still count or no?

mxddi3:

@snowflake0531 i don't see how this is incorrect. The question says the amount of copies (c) is (=) 400 per week (w). then they give `c=400w` so wouldn't this be correct? Also @Kayyy4oe welcome to questioncove. Could you specify what you mean when you say `is true of the graph of this scenario ? Aviable point on the graph is The values otw must be` ?

snowflake0531:

i said it's correct... wdym

Tranquility:

I have edited the question to reflect what it should actually read (without all the formatting errors).

jhonyy9:

Each Friday, the school prints 400 copies of the school newsletter. The equation c = 400w models the relationship between the number of weeks and the total number of copies of the newsletters printed. c = 400w c mean the total number of copies of the newsletters printed w mean the number of weeks given these points with coordinates and ask what can be one point of the graph ? A.(-2, -800) B.(2.5, 1,000) C.(6, 1,800) D.(8, 3,200) - so to get what is true you need to know that from this form of (x,y) x mean the weeks number and y the total number of copies printed so in this way x will be w in this formula and y will be c - to get the correct answer just you need check these coordinates substituted inside given equation and look what will result true for example the first one A. (-2,-800) substituted -800 = 400*(-2) -800 = -800 so yes this resulted true but first of all you need decid when you make something like copie can you make it in minus and resulted it in minus ? so i think NO ! B. (2.5 , 1000) substituted 1000 = 400*2,5 so what you think about this - can be true ? hope helped easy understandably

jhonyy9:

@AZ opinion pls - ty .

AZ:

You also have to keep in mind that weeks are whole numbers. You cannot have 1.5 weeks. It also has to be a positive number. You cannot have -3 weeks. That should help get you to your final answer :)